an investor invest 2500 in two mutual funds, he invest 5% in one and 2% in the other, both funds totaled an earning of $95, how much did each fund earn?
Answer by mananth(16949)

Part I 5.00% per annum ------------- Amount invested =x
Part II 2.00% per annum ------------ Amount invested = y
2500
Interest----- 95

Part I 5.00% per annum ---x
Part II 2.00% per annum ---y
Total investment
x + 1 y= 2500 -------------1
Interest on both investments
5.00% x + 2.00% y= 95
Multiply by 100
5 x + 2 y= 9500.00 --------2
Multiply (1) by -5
we get
-5 x -5 y= -12500.00
Add this to (2)
0 x -3 y= -3000
divide by -3
y = 1000
Part I 5.00% $ 1500
Part II 2.00% $ 1000

CHECK
1500 --------- 5.00% ------- 75.00
1000 ------------- 2.00% ------- 20.00
Total -------------------- 95.00

m.ananth@hotmail.ca